Solve for X and Y
2X + 7Y= -31
4X -4Y =-8

Answer by girl_from_moscow(7)

alrighty here we go )))
so first we will get rid of the x's
multiply the first equaition by (-2) because we need a -4x, because -4x+4x = 0
see?
so our first equation becomes -4x+-14=62
and our second equation stays the same.
-4x-14y=62
4x-4y =-8
add all the terms together and you will end up with:-18y=54
y=-3
what we now have to do in order to find x is to plug in the y into the ORIGINAL equation(it doesn't matter which one)
4x-4(-3)=-8
4x+12=-8
4x=-20
x=-5
so your final answer is y=-3 and x=-5 we used the elimination method